OpenStudy (anonymous):

Find the instantaneous rate of change of y with respect to x at an arbitrary value of x(v not)

OpenStudy (anonymous):

\[y=\frac{ 1 }{ x^2 }\] \[x_{o}=1\] \[x_{1}=\]

OpenStudy (anonymous):

\[x_{1}=2\]

zepdrix (zepdrix):

what is the x(v not)? Is that supposed to be a caret pointing down? like subscript? :) Use underscore for that. So we want to find y' evaluated at x_naut? Oh and they gave us a value for that? neato

OpenStudy (anonymous):

yea.

zepdrix (zepdrix):

Let's use rules of exponents to fix our function up a little bit.\[\Large y=\frac{1}{x^2}\qquad\to\qquad y=x^{-2}\]

zepdrix (zepdrix):

Have you learned about the `Power Rule` for derivatives?

OpenStudy (anonymous):

no, not yet.

zepdrix (zepdrix):

So we're using the `Limit Definition of the Derivative` to solve this?

OpenStudy (anonymous):

I am just starting derivatives, in the first section of the chapter, the ones with the slope and formula for tangent and secant

OpenStudy (anonymous):

yeah if u call it that

zepdrix (zepdrix):

So here is our thing we'll use:\[\Large y'=\lim_{h\to0}\frac{y(x+h)-y(x)}{x}\] I'm writing y in function notation, hope that's not too confusing :o

zepdrix (zepdrix):

AHHH typo, lemme fix that

zepdrix (zepdrix):

\[\Large y'=\lim_{h\to0}\frac{y(x+h)-y(x)}{h}\]

zepdrix (zepdrix):

You're probably more familiar with seeing f(x+h)-f(x) on top. Is it too confusing doing it this way?

OpenStudy (anonymous):

no its ok, but is it possible to use this one,\[y ^{'}=\lim_{x _{1} \rightarrow x _{o}}\frac{ f(x _{1})-f(x _{o}) }{ x _{1}-x _{o} }\]

OpenStudy (anonymous):

im not really accustomed to the h, but the teacher has showed us that eqtn

zepdrix (zepdrix):

Ah yes that form :) that would make more sense.

zepdrix (zepdrix):

mm one sec :3

OpenStudy (anonymous):

sure, no problem

zepdrix (zepdrix):

I don't think we want to use both initial x values for that form. The other limit formula, (the one you're used to using), is usually written as:\[\LARGE f'(x_o)=\lim_{x\to x_o}\frac{ f(x)-f(x _{o})}{x-x_o}\]Where we let the first x vary.

zepdrix (zepdrix):

Sorry I haven't used that limit formula very often lol :) maybe i'm being silly, lemme work it out on paper real quick.

OpenStudy (anonymous):

the first x meaning \[x _{o}\]?

zepdrix (zepdrix):

The first x meaning x. As in, leave it as a variable, plug it in and simplify the thing down. Then after we've gotten it into a nice form, we'll let "x approach x_o". (plug in x_o for our x's).

zepdrix (zepdrix):

It's really extra confusing that we're using y AND f(x) notation. Does this make sense on the left side of the equation if I write: \[\LARGE f'(x_o)=\lim_{x\to x_o}\frac{ f(x)-f(x _{o})}{x-x_o}\] Instead of y'? In which case we should write our original function as:\[\Large f(x)=\frac{1}{x^2}\]

zepdrix (zepdrix):

\[\Large \color{royalblue}{f(x)=\frac{1}{x^2}}\]\[\Large \color{orangered}{f(x_o)=\frac{1}{x_o^2}\qquad\to\qquad f(1)=\frac{1}{1^2}=1}\] We want to use these two pieces of information and plug them into our limit. \[\LARGE f'(x_o)=\lim_{x\to x_o}\frac{ \color{royalblue}{f(x)}-\color{orangered}{f(x _{o})}}{x-x_o}\]

zepdrix (zepdrix):

I color-coded them, maybe it will help match things up D:

zepdrix (zepdrix):

Understand how we're going to plug those pieces in? Or still way too confused on the limit formua that I used? :3

OpenStudy (anonymous):

it makes sense but the answer i got for this didnt match the answer key, no matter how i manipulated it, i was still alittle off from the book's answer.

zepdrix (zepdrix):

Plugging in those two pieces of information gives us this so far:\[\Large \lim_{x\to x_o}\frac{ \color{royalblue}{\dfrac{1}{x^2}}-\color{orangered}{1}}{x-1}\]

zepdrix (zepdrix):

Getting a common denominator up top gives us,\[\Large \lim_{x\to x_o}\frac{\left(\dfrac{1-x^2}{x^2}\right)}{x-1}\]

zepdrix (zepdrix):

Oh we should have a 1 in place of that x_o under our limit also.

zepdrix (zepdrix):

We can combine our denominators like so:\[\Large \lim_{x\to1}\dfrac{1-x^2}{x^2(x-1)}\]

OpenStudy (anonymous):

yes i get it, but that was not the way i did it, but ok

zepdrix (zepdrix):

This problem is a tad tricky, a lot of little steps to get there D:

zepdrix (zepdrix):

The numerator can be factored into the `difference of squares`:\[\Large \lim_{x\to1}\dfrac{(1-x)(1+x)}{x^2(x-1)}\] We'll factor out a negative 1 from the (1-x) term.\[\Large \lim_{x\to1}\dfrac{-(x-1)(1+x)}{x^2(x-1)}\] Which gives us a nice cancellation.

OpenStudy (anonymous):

k, that i understand also

zepdrix (zepdrix):

\[\Large \lim_{x\to1}\dfrac{-\cancel{(x-1)}(1+x)}{x^2\cancel{(x-1)}}\]From this form we can finally plug in x=1 (allowing x to approach 1).\[\Large \lim_{x\to1}\dfrac{-(1+x)}{x^2}\]

zepdrix (zepdrix):

What do you get?? :D

OpenStudy (anonymous):

ok, thats getting close

OpenStudy (anonymous):

im getting all the algebra ur doing

zepdrix (zepdrix):

You should be able to get your final answer by plugging x=1 directly in from that point.

OpenStudy (anonymous):

actually that was the solution that i got frist time i did it, then i checked with my teacher and he shrugged and said x naut should be included in your final answer. so therefore i did it again and i got stuck in factoring...

zepdrix (zepdrix):

That doesn't make any sense, why would x_o be included in your final answer if they provide a value for it.... that just seems silly -_-

OpenStudy (anonymous):

i dunno about that, cuz this is just c) of a problem, maybe reread the question i have up there? i have no idea

zepdrix (zepdrix):

If you're supposed to leave x_o in the equation, it would work out largely the same way.\[\Large f'(x_o)\quad=\quad\lim_{x\to x_o}\frac{\dfrac{1}{x^2}-\dfrac{1}{x_o^2}}{x-x_o}\]

zepdrix (zepdrix):

Common denominator up top gives us,\[\Large f'(x_o)\quad=\quad\lim_{x\to x_o}\frac{\left(\dfrac{x_o^2-x^2}{x^2x_o^2}\right)}{x-x_o}\]

zepdrix (zepdrix):

That step is a little tricky :o understand how to combine those fractions?

OpenStudy (anonymous):

omgAAAAD, i was typing so much and the page broke....

OpenStudy (anonymous):